Calculate the amount of heat (joules) required to convert 200 g of ice cubes (00 C) to gas at 100 0
1
Expert's answer
2020-12-18T08:43:10-0500

1. Determine the amount of neat absorbed by the ice during melting:

"Q_f = 200 \times 333 = 66600 \;J"

2. Use the specific heat equation to determine the amount of heat absorbed by the water until its reaches the boiling point.

Q_w = mCΔT

"= 200 \times 4.184 \times 100 = 83680 \;J"

3. Determine the amount of heat absorbed by the water during vaporization.

"Q_v = 200 \times 2260 = 452000 \:J"

4. Calculate the total energy absorbed during the entire process.

"Q_T = Q_f+Q_w+Q_v \\

Q_T = 66600+83680+452000=602280 \;J = 602.28 \;kJ"

Answer: 602.28 kJ